NO2(g)  ⇌ N2O4(g)             Kp = 11.3 at 25oC

(a) Calculate G if the P(NO2) = 0.29 atm, and P(N2O4) = 1.6 atm.

(b) Predict the direction in which the reaction will shift to reach equilibrium. Explain how you determined this.

(c) Once equilibrium is reached are there more reactants or products present. Explain how you determined this.
Answer #1

kp=113 at es c. NO(g) - HQOA (8) iTo 95c : 298k. AG°E-RT nk. AG = AGO + RT ln Q. PLM22] = 1-6 afm. a) Given- PLNO2] = 0.29cm

Hope you will like my answer.

If any doubt please comment.